Please answer the following question with steps and explanation (I am very confused on this problem).
2 = x-2
___ ____
x-3 6
Thanks :)

Answer by chessace(471)

2/(x-3) = (x-2)/6
Cross mulltiply to get
12 = (x-2)*(x-3)
Flip and expand right hand side:
(x^2 -5x +6) -12 =0
Regroup:
x^2 - 5x - 6 = 0
Factor:
(x-6)(x+1) = 0
x = 6, -1.